# Cooperation Does Matter: Exploring Multi-Order Bilateral Relations for Audio-Visual Segmentation

### 11 Dec 2023 v1

@@ 3.3 Mining Temporal Relationships
Adaptive Inter-frame Consistency Loss. Temporal always implies a bilateral relationship in nature. For instance, in video clips, one can deduce the scenario of the current frame based on the past frame. Similarly, it is also feasible to predict the future frame based on the current frame. This interactive dynamic among frames can be construed as a type of temporal entanglement. To take advantage of this potential temporal entanglement, we introduce an Adaptive Inter-Frame Consistency Loss for the AVS task. The similarity score for each successive frame can be calculated using the following formula:
@@
where $O_{t}^{mask}$ refers to the predicted masks at frame $t$, and $\text{cos}(\cdot)$ symbolizes the cosine similarity function. The term $\mathcal{S}{t:t+1}$ represents the similarity score between frames $t$ and $t+1$. Upon examination in Fig. 4, it is evident that a significant similarity exists between distinct frames. Therefore, to leverage this prior information, we propose an adaptive inter-frame consistency loss, formulated as follows:
$\mathcal{L}_{ada}=\sum_{t=1}^{T-1}\exp(-\mathcal{S}_{t:t+1})(1-\mathcal{S}_{t:t+1}),$ (7)
where $\exp(-\mathcal{S}_{t:t+1})$ represents adaptive weight. When the disparity between adjacent frames is substantial, the adaptive weight item will be minimal, aligning with intuition.
@@ 3.4 Training and Inference
Overall Training Loss. The comprehensive training loss comprises three components: mask loss, classification loss, and adaptive inter-frame consistency loss, as previously discussed. The classification loss is formulated by a cross-entropy loss, denoted as $\mathcal{L}_{cls}=\mathcal{L}_{ce}$. Mask loss integrates the binary cross-entropy loss and the dice loss [32], and is depicted as $\mathcal{L}_{mask}=\mathcal{L}_{bce}+\mathcal{L}_{dice}$. Considering that in the AVS task, the ratio of segmented objects to the total image area is relatively small, employing dice loss allows the model to better concentrate on the foreground and minimizes distraction from the background. The overall training loss is expressed as follows:
@@
Semantic Inference. After obtaining the predicted embeddings $O^{cls}\in\mathbb{R}^{T\times N_{q}\times(K+1)}$ and binary masks $O^{mask}\in\mathbb{R}^{T\times N_{q}\times H\times W}$, where $K$ represents the total number of object classes and $N_{q}$ is the number of object queries, we employ the identical post-processing as in [10] to yield the final semantic segmentation outputs. Speciafilly, we first calculated the output mask with classes $O=O^{cls}\times O^{mask}\in\mathbb{R}^{T\times(K+1)\times H\times W}$ Then, we execute $\arg\max$ and discard the no object class to obtain the ultimate results.
@@
(+2.6) (+2.0) (+1.2) (+1.2)
Table 1: Quantitative comparison results of different methods on AVSBench-object (Single-source, S4; Multi-source, MS3). We use the same backbones (ResNet-50 and PVTv2) to demonstrate that our method outperforms other methods significantly.
@@ 4.1 AVSBench Datasets
Table 2: Quantitative comparison results of different methods on AVSBench-semantic (AVSS).
@@
$w/o$ Siam-Backbone 80.6 88.7 31.9 35.7
$w/o$ Siam-Fusion 81.1 89.9 33.1 36.7
@@ 4.2 Experimental Setup
Implementation Details. For fair comparison, we adopt the ImageNet pre-trained ResNet-50 [16] and Pyramid Vision Transformer (PVT-v2) [36] as the visual siam-encoders. All input frames are resized to $224\times 224$. In terms of the proposal generator, we leverage the Semantic-SAM [22] to obtain class-agnostic masks. As for audio input, we take the Vggish encoder [17] pre-trained on AudioSet [14] to extract audio features. Following [10], the Multi-Scale Deformable Attention Transformer (MSDeformAttn) is our default pixel decoder. Besides, we adopt the standard Transformer Decoder, with $L=3$ (i.e., a total of 9 layers) and $N_{q}=100$ as the default. The hyperparameters is set as $\lambda_{cls}=2$ and $\lambda_{mask}=5$. For the inter-frame consistency loss, we set $\lambda_{ada}=10$ for the AVSBench-object dataset, $\lambda_{ada}=5$ for the AVSBench-semantic dataset due to longer frames per video. We calculate the inter-frame consistency loss only in the intermediate Transformer decoder layer to prevent an over-dependence on temporal information. All models are trained using the Adam optimizer with a learning rate of $41e-$ 4 and weight decay of $0.05$. We train the S4 and AVSS subsets for 90k iterations and MS3 for 20k iterations with a batch size of 8 on a single Tesla A100 40GB GPU.
Metrics. Folloing [45, 13, 31], we adopt two metrics to verify the effectiveness, namely, Jaccard index $\mathcal{J}$ and F-score $\mathcal{F}$. $\mathcal{J}$ computes the intersection over union (IoU) of the predicted segmentation and the ground truth mask. $\mathcal{F}$ considers both precision and recall, which is represented as $\mathcal{F}_{\beta}=\frac{(1+\beta^{2})\times\text{presion}\times\text{recall}}{\beta^{2}\times\text{precision}+\text{recall}}$, where $\beta^{2}$ is set at $0.3$. In our experiment, we use $\mathcal{M_{J}}$ and $\mathcal{M_{F}}$ to denote the mean metrics across the entire dataset.
@@ 4.3 Main Results
We conduct experiments on AVSBench-object and AVSBench-semantic datasets. As AVS is an emerging proposed problem recently introduced by [45], we compare our COMBO with some state-of-the-art methods from other related tasks, such as sound source localization (SSL) [5, 34], video object segmentation (VOS) [28, 12, 39], and salient object detection (SOD) [29, 41], all of which provide a comparative benchmark for our experiments. As evidenced in Tab. 1, COMBO demonstrates a substantial performance gap ($+9.8$ mIoU in S4; $+16.3$ mIoU in MS3) over other related methods, principally attributable to variances in setting specific task scenarios. We also compare our method against some recent state-of-the-art methods [45, 15, 23, 30, 31, 26, 13] that have been explicitly designed for audio-visual segmentation settings. On the AVSBench-object dataset, COMBO-R50 outperforms the current best performance by achieving 3.7 mIoU and 4.8 F-score improvements for S4 subset and 2.7 mIoU and 1.3 F-score improvements for MS3, while COMBO-PVT surpasses the top-performing method by 2.6 mIoU and 2.0 F-score for S4 and 1.2 mIoU and 1.2 F-score for MS3.
Besides, we compare the AVSBench-semantic dataset as displayed in Tab. 2, which presents a more challenging setting. Both COMBO-R50 and COMBO-PVT achieve significant results, with 8.4 and 5.4 mIoU improvements and significant F-score enhancements of 8.0 and 4.1, respectively. These experiments confirm that our COMBO model surpasses the performance of existing state-of-the-art methods across all sub-tasks, consequently setting a new benchmark for audio-visual segmentation.
@@ 4.4 Abaltion Study
Component analysis of COMBO. To validate the impact of our proposed method, we separately eliminate the Siam-Encoder Module (SEM), Bilateral-Fusion Module (BFM), and Inter-Frame Similarity Loss. As demonstrated in Tab. 3, the results indicate that COMBO has demonstrated superior influence on SEM with 1.4 mIoU improvement, particularly on the AVSS subset. Concurrently, our findings indicate that BFM is of substantial significance, demonstrating a performance enhancement of 0.6 mIoU over the S4 subset. More analysis will be discussed later. In addition, we examine the effect of inter-frame consistency loss. The results reveal that our loss function contributes to performance improvements with a 0.7 mIou increase on the S4 subset. More details will be provided subsequently.
Effects of Siam-Encoder Module (SEM). We first examine the significance of the Siam-Encoder Module. Two designs are compared in Tab. 4: one with share weights and another with separate weights. When comparing shared and separate parameters, it is found that separable parameters can attain 0.2 mIoU on the S4 subset, but it is indeed more costly than the others.
Effects of Bilateral-Fusion Module (BFM). Moreover, we investigate the influence of the BFM. We compared our modules using different variants, which include without any fusion, video-only fusion (inject audio feature into visual), audio-only fusion (inject visual feature to audio), and our bilateral fusion. As demonstrated in Tab. 5, bilateral fusion achieves a performance improvement of 0.6 mIoU compared to the model without any fusion. It is worth noting that the transformer decoder inherently consists of a cross-attention function, potentially serving as a fusion process. Our model also realizes a performance improvement of 0.4 mIoU compared to only audio or visual fusion.
Effects of Audio Queries. As demonstrated in Tab. 6, All denotes the exclusive use of fused audio queries, and Add signifies the combination of fused audio queries and learnable queries. We first expand the audio features to the exact dimensions as learnable queries, then compare the experiments with the exclusive use of fused audio queries and the combination of fused audio queries and learnable queries. The results show that query with Add have 0.8 and 2.6 mIoU improvements over query with All on S4 and AVSS subsets.
@@ 4.5 Qualitative Analysis
We provide a qualitative comparison between our proposed method and [46, 13] on AVSBench-object and AVSBench-semantic datasets. As depicted in Fig.5, our proposed method, COMBO, exhibits superior audio-temporal and spatial localization quality, leading to better visualization and segmentation performance. For instance, in the case of the middle samples, our model accurately segments the singing man despite the presence of other sounds. Moreover, our method achieves more precise segmentation for background noise handling and provides a richer detail of the foreground in other examples.
@@ 5 Conclusion
We introduce a novel audio-visual transformer framework, termed COMBO, that archives state-of-the-art performance on AVSBench-object and AVSBench-semantic datasets. Contrary to previous methodologies that only factor in modality or temporal relations individually, our proposed method explores multi-order bilateral relations for the first time, combining pixel entanglement, modality entanglement, and temporal entanglement. For these three kinds of entanglement, we propose Siam-Encoder Module (SEM), Bilateral-Fusion Module (BFM), and inter-frame consistency loss, respectively. Extensive experimental results verify the effectiveness of our proposed framework. We hope that our work will inspire further research in this significant and worthwhile field.
@@ B.2 Proposal Generator
As shown in Fig. II, we provide a more detailed explanation of the proposal generator proposed in COMBO. Initially, we obtain class-agnostic masks denoted as $c\in\mathbb{R}^{K\times H\times W}$ from the input frame, using a pre-existing foundation model [22], where $K$ symbolizes the number of potential targets. Subsequently, a Maskige generator, which is part of the proposal generator, is introduced to convert the class-agnostic masks $c\in\mathbb{R}^{K\times H\times W}$ into Maskige, denoted as $m\in\mathbb{R}^{3\times H\times W}$ without the need for additional training. Particularly, as $K$ is dynamic and fluctuates according to input frames, we amplify the quantity of class-agnostic masks to $N$ using zero masks, thereby deriving a series of binary masks $c\in\mathbb{R}^{N\times H\times W}$, where $N$ is a predetermined number and $N\geq K$. Next, considering that $c$ denotes a series of binary masks that are challenging to incorporate into visual features, we utilize a random color encoding function $\mathcal{X}(\cdot):\mathbb{R}^{N\times H\times W}\rightarrow\mathbb{R}^{3\times H\times W}$ to convert the binary masks $c\in\mathbb{R}^{N\times H\times W}$ into Maskige $m\in\mathbb{R}^{3\times H\times W}$. Formally, we define $\mathcal{X}(c)=cA$, where $A\in\mathbb{R}^{N\times 3}$. To facilitate the proposal generator offline, the value of $A$ is manually set appropriately. Specifically, we set $N=100$, a value considerably larger than $K$. To enhance the distinctiveness between various targets, as illustrated in Fig. III, we use the first 100 color mappings of the color mapping relationship in ADE20K dataset [44] as the parameters of matrix $A$. Further visualizations on Maskiges are available in Sec. D.

### 07 Apr 2024 v2

@@ 3.3 Mining Temporal Relationships
Adaptive Inter-frame Consistency Loss. Temporal always implies a bilateral relationship in nature. For instance, in video clips, one can deduce the scenario of the current frame based on the past frame. Similarly, it is also feasible to predict the future frame based on the current frame. This interactive relations among frames can be construed as a type of temporal entanglement. To take advantage of this potential temporal entanglement, we introduce an adaptive inter-frame consistency loss for AVS. The similarity score for each successive frame can be calculated as follows:
@@
where $O_{t}^{mask}$ refers to the predicted masks at frame $t$, and $\text{cos}(\cdot)$ symbolizes the cosine similarity function. The term $\mathcal{S}_{t:t+1}$ represents the similarity score between frames $t$ and $t+1$. As illustrated in Fig. 4, it is evident that a significant similarity exists between distinct frames. Therefore, to leverage this prior information, we propose an adaptive inter-frame consistency loss, formulated as follows:
$\mathcal{L}_{ada}=\sum_{t=1}^{T-1}\exp(\mathcal{S}_{t:t+1}-1)(1-\mathcal{S}_{t:t+1}),$ (7)
where $\exp(\mathcal{S}_{t:t+1}-1)$ represents adaptive weight. When the disparity between adjacent frames is substantial, the adaptive weight item is minimal, aligning with intuition.
@@ 3.4 Training and Inference
Overall Training Loss. The comprehensive training loss comprises three components: classification loss, mask loss, and adaptive inter-frame consistency loss, as previously discussed. The classification loss is formulated by a cross-entropy loss, denoted as $\mathcal{L}_{cls}=\mathcal{L}_{ce}$. The mask loss integrates the binary cross-entropy loss and the dice loss [32], and is depicted as $\mathcal{L}_{mask}=\mathcal{L}_{bce}+\mathcal{L}_{dice}$. Considering that in the AVS task, the ratio of segmented objects to the total image area is relatively small, employing dice loss allows the model to better concentrate on the foreground and minimizes distraction from the background. The overall training loss is expressed as follows:
@@
Semantic Inference. After obtaining the predicted embeddings $O^{cls}\in\mathbb{R}^{T\times N_{q}\times(K_{c}+1)}$ and binary masks $O^{mask}\in\mathbb{R}^{T\times N_{q}\times H\times W}$, where $K_{c}$ represents the total number of object classes and $N_{q}$ is the number of object queries, we employ the identical post-processing as in [10] to yield the final semantic segmentation outputs. Specifically, we first calculate the output mask with classes $O=O^{cls}\times O^{mask}\in\mathbb{R}^{T\times(K_{c}+1)\times H\times W}$. Then, we execute $\arg\max$ and discard the no object class to obtain the ultimate results.
@@
(+2.6) (+2.0) (+0.2) (+1.2)
Table 1: Quantitative comparison results of different methods on AVSBench-object (Single-source, S4; Multi-source, MS3). We use the same backbones (ResNet-50 and PVT-v2) to demonstrate that our method outperforms other methods significantly.
@@ 4.1 AVSBench Datasets
Table 2: Quantitative comparison results on AVSBench-semantic.
@@
$w/o$ Siam-Encoder 80.6 88.7 31.9 35.7
$w/o$ Bilateral-Fusion 81.1 89.9 33.1 36.7
@@ 4.2 Experimental Setup
Implementation Details. For fair comparison, we adopt the ImageNet pre-trained ResNet-50 [16] and Pyramid Vision Transformer (PVT-v2) [36] as the visual siam-encoders. All input frames are resized to $224\times 224$. In terms of the proposal generator, we leverage the Semantic-SAM [22] to obtain class-agnostic masks. As for audio input, we take the Vggish encoder [17] pre-trained on AudioSet [14] to extract audio features. Following [10], the Multi-Scale Deformable Attention Transformer (MSDeformAttn) is our default pixel decoder. Besides, we adopt the standard transformer decoder, with $L=3$ (i.e., a total of 9 layers) and $N_{q}=100$ as the default. The hyperparameters are set as $\lambda_{cls}=2$ and $\lambda_{mask}=5$. For the inter-frame consistency loss, we set $\lambda_{ada}=10$ for the AVSBench-object dataset, $\lambda_{ada}=5$ for the AVSBench-semantic dataset due to longer frames per video. We calculate the inter-frame consistency loss only in the intermediate transformer decoder layer to prevent an over-dependence on temporal information. All models are trained using the Adam optimizer with a learning rate of $41e-$ 4 and weight decay of $0.05$. We train the S4 and AVSS subsets for 90k and MS3 for 20k iterations with a batch size of 8 on a single A100 40GB GPU.
Metrics. Folloing [45], we adopt two metrics to verify the effectiveness, namely, Jaccard index $\mathcal{J}$ and F-score $\mathcal{F}$. $\mathcal{J}$ computes the intersection over union (IoU) of the predicted segmentation and the ground truth mask. $\mathcal{F}$ considers both precision and recall, which is represented as $\mathcal{F}_{\beta}=\frac{(1+\beta^{2})\times\text{presion}\times\text{recall}}{\beta^{2}\times\text{precision}+\text{recall}}$, where $\beta^{2}$ is set at $0.3$. In our experiment, we use $\mathcal{M_{J}}$ and $\mathcal{M_{F}}$ to denote the mean metrics across the entire dataset.
@@ 4.3 Main Results
We conduct experiments on AVSBench-object and AVSBench-semantic datasets. As AVS is an emerging proposed problem recently introduced by [45], we compare our COMBO with some state-of-the-art methods from other related tasks, such as sound source localization (SSL) [5, 34], video object segmentation (VOS) [28, 12, 39], and salient object detection (SOD) [29, 41], all of which provide a comparative benchmark for our experiments. As evidenced in Tab. 1, COMBO demonstrates a substantial performance gap ($+9.8$ mIoU in S4; $+16.3$ mIoU in MS3) over other related methods, principally attributable to variances in setting specific task scenarios. We also compare our method against some recent state-of-the-art methods [45, 15, 23, 30, 31, 26, 13] that have been explicitly designed for audio-visual segmentation settings. On the AVSBench-object dataset, COMBO-R50 outperforms the current best performance by achieving 3.7 mIoU and 4.8 F-score improvements for S4 subset and 2.7 mIoU and 1.3 F-score improvements for MS3, while COMBO-PVT surpasses the top-performing method by 2.6 mIoU and 2.0 F-score for S4 and 0.2 mIoU and 1.2 F-score for MS3.
Besides, we compare the AVSBench-semantic dataset as displayed in Tab. 2, which presents a more challenging setting. Both COMBO-R50 and COMBO-PVT achieve significant results, with 8.4 and 5.4 mIoU improvements and significant F-score enhancements of 8.0 and 4.1, respectively. These experiments confirm that our COMBO model surpasses existing state-of-the-art methods across all sub-tasks, consequently setting a new benchmark for AVS.
@@ 4.4 Abaltion Study
Component analysis of COMBO. To validate the impact of our proposed method, we separately eliminate the Siam-Encoder Module (SEM), Bilateral-Fusion Module (BFM), and adaptive inter-frame consistency loss. As demonstrated in Tab. 3, the results indicate that COMBO has demonstrated superior influence on SEM with 1.4 mIoU improvement, particularly on the AVSS subset. Concurrently, our findings indicate that BFM is of substantial significance, demonstrating a performance enhancement of 0.6 mIoU over the S4 subset. More analysis are discussed later. In addition, we examine the effect of inter-frame consistency loss. The results reveal that our loss function contributes to performance improvements with 0.7 mIoU on the S4 subset. More details are provided subsequently.
Effects of Siam-Encoder Module (SEM). We first examine the significance of our SEM. Two designs are compared in Tab. 4: one with share weights and another with separate weights. When comparing shared and separate parameters, it is found that separable parameters can attain 0.2 mIoU on the S4 subset, but it is indeed more costly than the others.
Effects of Bilateral-Fusion Module (BFM). Moreover, we investigate the influence of our BFM. We compared our modules using different variants, which include without any fusion, video-only fusion (inject audio feature into visual), audio-only fusion (inject visual feature to audio), and our bilateral fusion. As demonstrated in Tab. 5, bilateral fusion achieves a performance improvement of 0.6 mIoU compared to the model without any fusion. It is worth noting that the transformer decoder inherently consists of a cross-attention function, potentially serving as a fusion process. Our model also realizes a performance improvement of 0.4 mIoU compared to only audio or visual fusion.
Effects of Audio Queries. As demonstrated in Tab. 6, All denotes the exclusive use of fused audio queries, and Add signifies the combination of fused audio queries and learnable queries. We first expand the audio features to the exact dimensions as learnable queries, then compare the experiments with the exclusive use of fused audio queries and the combination of fused audio queries and learnable queries. The results show that queries with Add have 0.8 and 2.6 mIoU improvements over All on S4 and AVSS subsets.
@@ 4.5 Qualitative Analysis
We provide a qualitative comparison between AVSBench [46], AVSegformer [13] and our proposed method on AVSBench-object and AVSBench-semantic datasets. As depicted in Fig. 5, our method, COMBO, exhibits superior audio-temporal and spatial localization quality, leading to better visualization and segmentation performance. For instance, in the case of the middle samples, our model accurately segments the singing man despite the presence of other sounds. Moreover, our method achieves more precise segmentation for background noise handling and provides richer details of the foreground in other examples.
@@ 5 Conclusion
We introduce a novel audio-visual transformer framework, termed COMBO, that archives state-of-the-art performance on AVSBench-object and AVSBench-semantic datasets. Contrary to previous methodologies that only factor in modality or temporal relations individually, our method explores multi-order bilateral relations for the first time, combining pixel entanglement, modality entanglement, and temporal entanglement. For these three kinds of entanglement, we propose Siam-Encoder Module (SEM), Bilateral-Fusion Module (BFM), and adaptive inter-frame consistency loss, respectively. Extensive experimental results verify the effectiveness of our proposed framework. We hope that our work will inspire further research in this significant and worthwhile field.
Acknowledgements
This research was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB0500103), the National Natural Science Foundations of China (Grants No. 62076242, 62376267), the Pre-Research Project on Civil Aerospace Technologies (No. D030312), the National Defense Basic Scientific Research Program of China (No. JCKY2021203B063) and the innoHK project.
@@ B.2 Proposal Generator
As shown in Fig. II, we provide a more detailed explanation of the proposal generator proposed in COMBO. Initially, we obtain class-agnostic masks denoted as $c\in\mathbb{R}^{K\times H\times W}$ from the input frame, using a pre-existing foundation model [22], where $K$ denotes the number of potential targets. Subsequently, a Maskige generator, which is part of the proposal generator, is introduced to convert the class-agnostic masks $c\in\mathbb{R}^{K\times H\times W}$ into Maskige, denoted as $m\in\mathbb{R}^{3\times H\times W}$ without the need for additional training. Particularly, as $K$ is dynamic and fluctuates according to input frames, we amplify the quantity of class-agnostic masks to $N$ using zero masks, thereby deriving a series of binary masks $c\in\mathbb{R}^{N\times H\times W}$, where $N$ is a predetermined number and $N\geq K$. Next, considering that $c$ denotes a series of binary masks that are challenging to incorporate into visual features, we utilize a random color encoding function $\mathcal{X}(\cdot):\mathbb{R}^{N\times H\times W}\rightarrow\mathbb{R}^{3\times H\times W}$ to convert the binary masks $c\in\mathbb{R}^{N\times H\times W}$ into Maskige $m\in\mathbb{R}^{3\times H\times W}$. Formally, we define $\mathcal{X}(c)=cA$, where $A\in\mathbb{R}^{N\times 3}$. To facilitate the proposal generator offline, the value of $A$ is manually set appropriately. Specifically, we set $N=100$, a value considerably larger than $K$. To enhance the distinctiveness between various targets, as illustrated in Fig. III, we use the first 100 color mappings of the color mapping relationship in ADE20K dataset [44] as the parameters of matrix $A$. Further visualizations on Maskiges are available in Sec. D.